OpenStudy (anonymous):

Someone help me please :c What is the simplified form of the expression? c^9 d^-7 / c^14 d^-10

OpenStudy (anonymous):

you can change your equation to \[c^9*d^{10}/c^{14}*d^7\]

OpenStudy (anonymous):

Sorry for the confusion I don't know how to make the little numbers

OpenStudy (anonymous):

you can then subtract exponents on the like terms so c^9-c^14 and then d^10-d^7

OpenStudy (anonymous):

so you should get c^-5 * d^3

OpenStudy (anonymous):

c^9-c^14=c-5 and d^10-d^7= d3 ?

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

c^-5 * d^3

OpenStudy (anonymous):

which simplified is \[d^3/c^5\]

OpenStudy (anonymous):

wow, that's all you have to do? so simple x.x

OpenStudy (anonymous):

I'm going to save this conversation for any other questions I get that look like this. thanks so much

OpenStudy (anonymous):

and if you want to check it you can always plug in numbers into the original and the simplified equation to see if you get the same answer

OpenStudy (anonymous):

Oh and the exponent can be made by typing them into the equation generator at the bottom of the post box and use {} around the number if they are more than one digit

OpenStudy (anonymous):

Ohh, okay, thanks I'll use that next time too =D